OpenStudy (anonymous):

solve for x x/2+x-5=15

OpenStudy (anonymous):

(1/2) * x + x - 5 = 15. so you actually (1/2)*x + (2/2)*x -5 = 15

OpenStudy (anonymous):

combine the x terms, the isolate x

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

np

OpenStudy (anonymous):

\[x/2 + x - 5 = 15\]